OpenStudy (anonymous):

What are the solutions to the following system of equations? Select the correct answer below. y = x2 − 4x + 3 2x + y = 6

OpenStudy (anonymous):

@ganeshie8

OpenStudy (hybrik):

To do this, you have to order the equation: y = x2 − 4x + 3 y = -2x + 6 Subtract Y x2-6x=-3 to find X, you must solve by substitution. (-2x+6) = x2 − 4x + 3 -2x+6=x2-4x+3 x2-2x-3=0 Then factor: 1x -3 1x 1 (x-3)(x+1) x = 3,-1 (3, y), (-1, y) to find this you plug in X to any equation: y = -2(3) + 6 y=-6+6 y=0 y = -2(-1) + 6 y=2+6 y=8 > (3,0), (-1,8) <

OpenStudy (anonymous):

Thanks. You are amazing xoxox